OpenStudy (anonymous):

Can anyone help with Converting functions into standard form? will give medals for best response

OpenStudy (anonymous):

Y=2(x-4)(x+3)

OpenStudy (anonymous):

Y-2=-3(x+4)^2

OpenStudy (anonymous):

and y=-x^2+5x+36

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@myininaya and @Hero can you please help me :)

hero (hero):

Standard Form of a Quadratic Equation: \[y = ax^2 + bx + c\] Standard Form of a Quadratic Function: \[F(x) = a(x - h)^2 + k\]

OpenStudy (anonymous):

i think its the second formula